# Understanding Self-Admitted Technical Debt in Test Code: An Empirical Study

### v1

@@ 5.3.1. Production-originated Issues
Failures: This subcategory of SATD indicates the causes of test failures, which are related to either environments/settings or inputs in the production code. 15 cases fell into this category, with 11 related to environments/settings and 4 related to specific inputs. The most two frequently words are fails and enabled, which is related to test execution status. Bavota et al. (Bavota and Russo, 2016) also introduced examples of such failures, but their “Test debt” category does not contain any subcategories. Similarly, the study by Kashiwa et al. (Kashiwa et al., 2022) also presents a “Failure” category for test-related SATD, but no further categorization was given. We separate the “failure”-related SATD into two types:
  • Indicates the failure-triggering environments/settings: This type of SATD is used to notify developers that a test will fail with a specific environment or configuration. In many cases, developers comment out the assertions that fail or disable the whole test method. Snippet 8 shows an example, including an access-related failure occurring only on Linux and Windows. [^11] https://github.com/apache/commons-io/blob/290d72eda9152d1e11b79d48453908ff3f6b9897/src/test/java/org/apache/commons/io/FileUtilsTest.java#L1730 This test method is enabled to run only on Mac OS, using “@EnabledOnOs” annnotation. [^12] https://junit.org/junit5/docs/5.2.0/api/org/junit/jupiter/api/condition/EnabledOnOs.html
@@
  • Indicates the failure-triggering inputs: This type of SATD notifies developers that the test will fail when the production code receives a specific input, which is illustrated in Snippet 9. In the example, the test is commented out due to a failure caused by a specific file. [^13] https://github.com/argouml-tigris-org/argouml/blob/6b6db0242a40f80655cbfdddcca246afe23df20c/src/argouml-core-model-mdr/tests/org/argouml/model/mdr/TestReadCompressedFilesAndHref.java#L73
@@
    • Indicates the existence of defects in production code: This type of SATD is used to inform developers of defects present in the production code. We identified 42 instances of this type, making it the most prevalent on-hold task. Snippet 10 shows an example of this category, where the test is currently disabled until the defect in the YARN project is fixed. [^14] https://github.com/apache/hadoop/blob/bd8b77f398f626bb7791783192ee7a5dfaeec760/hadoop-mapreduce-project/hadoop-mapreduce-client/hadoop-mapreduce-client-app/src/test/java/org/apache/hadoop/mapreduce/v2/app/TestStagingCleanup.java#L224-L227
@@
    • Indicates the unimplemented production code: This type of SATD indicates that a developer is waiting for the implementation of certain production code. An example can be seen in Snippet 11, in which a method call is commented out because a function for a property is not yet implemented. [^15] https://github.com/apache/poi/blob/ae2f0945cd2ab37260e46ab46c54b8f68a131aea/poi/src/test/java/org/apache/poi/hpsf/basic/TestWriteWellKnown.java#L230
@@
    • Specifies tasks to execute for debugging: This type of SATD informs developers of the specific task to execute for debugging. As illustrated in Snippet 12, the SATD comment asks other developers to investigate why an item in a JSON file was not generated. [^16] https://github.com/apache/camel/blob/66736471db8ddc22e50cc31c87d34b072455b488/components/camel-aws/camel-aws-xray/src/test/java/org/apache/camel/component/aws/xray/ErrorTest.java#L43
@@ 5.3.2. Test-originated Issues
        Indicates workaround for impossible/difficult implementations: This type of SATD is used to inform developers about the workaround for code either impossible or difficult to implement. This is the most frequent SATD type (53 instances) regarding Test Design and Implementation. An example can be seen in Snippet 15, in which a developer attempted to retrieve the value of a property, but the current implementation of the production code prevents this due to encapsulation. Consequently, the developer resorted to using reflection to obtain the value. The comment clarifies the rationale behind this implementation choice. [^19] https://github.com/quarkusio/quarkus/blob/a91a36c533676a5d35ddedbef1717392c9191360/integration-tests/kafka-streams/src/test/java/io/quarkus/it/kafka/streams/KafkaStreamsPropertiesTest.java#L26
@@
        Indicates non-optimal way of test implementation: This type of SATD is used to highlight non-optimal implementations. Snippet 16 provides an example of a such case, where a comment points out the use of a method or approach that is suboptimal. In this instance, the SATD comment indicates that directly using DataFormatReifier is not the best practice and suggests that a better approach should be considered. [^20] https://github.com/apache/camel/blob/66736471db8ddc22e50cc31c87d34b072455b488/components/camel-fhir/camel-fhir-component/src/test/java/org/apache/camel/component/fhir/dataformat/spring/FhirDataformatConfigSpringTest.java#L90
@@
        Provides an alternative way to implement/design the test: This type of SATD is used to propose a different way to implement the test. Snippet 17 provides an example of this type of SATD, where the comment suggests using “interrupts” instead of notifyAll() to target waiting threads more effectively. [^21] https://github.com/apache/cassandra/blob/6b134265620d6b39f9771d92edd29abdfd27de6a/test/simulator/main/org/apache/cassandra/simulator/systems/InterceptingMonitors.java#L363
@@
        Doubts on specific test case design: This type of SATD raises questions about purposes, validity or effectiveness of test cases. Snippet 18 provides an example where a developer questioned the purpose of the test case and subsequently commented out the entire test method. [^22] https://github.com/apache/bookkeeper/blob/cce4b6461691466c663f2cb4d00dd4d73dd9071e/bookkeeper-server/src/test/java/org/apache/bookkeeper/bookie/datainteg/DataIntegrityCheckTest.java#L410
@@
        Doubts on test implementation: This type of SATD appears when developers are unclear about whether the implementations are optimal or not. Snippet 19 provides an example, in which the comment raises questions about the necessity of the Thread.sleep() call, suggesting that a further review is needed. [^23] https://github.com/apache/camel/blob/66736471db8ddc22e50cc31c87d34b072455b488/components/camel-github/src/test/java/org/apache/camel/component/github/consumer/PullRequestCommentConsumerTest.java#L55
@@
        Indicates the non-functional issues of tests: This type of SATD is used to inform developers about non-functional issues in the test code, such as performance, scalability, or resource usage. Snippet 20 provides an example, where the comment indicates that the test takes an excessive amount of time to execute, potentially impacting overall test efficiency. [^24] https://github.com/cloudfoundry/uaa/blob/5b74878c0297043860bb88434ee1123934acfe19/server/src/test/java/org/cloudfoundry/identity/uaa/util/CachingPasswordEncoderTest.java#L125
@@
        Test Maintenance: This subcategory of SATD is related to test code maintenance. These SATD instances indicate the need for updates in future versions, refactoring, fixing flaky tests, and adding documentation. Frequent terms such as remove and version indicate that this category captures SATD related to test code management, including cleanup and version updates. This subcategory does not exist in Bavota et al.’s work (Bavota and Russo, 2016), but “Asks for test updates in response to future/latest software versions” partially matches the “Future work” subcategory in Kashiwa et al.’s study (Kashiwa et al., 2022). We identified 48 cases in this category, with the most frequent subcategory being “Asks for test updates in response to future/lastest software versions,” represented by 17 instances. The details of each subcategory are described below.
          Asks for test updates in response to future/lastest software versions: This type of SATD is used to inform developers that the test needs to be updated in a future version of production code. Snippet 21 provides an example of this subcategory. The comment suggests the removal of the reference to SecurityManager in this test case when the project updates the version of the programming language Groovy, as the class is discontinued in the new version. [^25] https://github.com/apache/groovy/blob/f6221ee780bfb2f84fb197da2c13387c4e93a019/src/test/org/codehaus/groovy/reflection/SecurityTest.java#L301
          Example of SATD categorized into “Asks for test updates in response to future/lastest software versions”
@@
          Asks for test code restructuring/refactoring: This type of SATD is used to notify developers that the test code needs to be restructured or refactored. Snippet 22 shows an example, where the SATD comment suggests making the contract more explicit or extracting common code to improve readability and maintainability. [^26] https://github.com/apache/flink/blob/eaffd227d853e0cdef03f1af5016e00f950930a9/flink-state-backends/flink-statebackend-changelog/src/test/java/org/apache/flink/state/changelog/ChangelogStateDiscardTest.java#L376 When looking into 9 cases, most of the SATDs request structural refactoring, including five instances of Move Method/Class, two of Extract Method, one of Rename Method, and one involving data structure changes (e.g., Introduce Parameter Object). We also found that only one case co-occurred with a test smell (specifically, “Assertion Roulette”), reinforcing our finding that Test SATD and test smells are largely independent issues.
@@
          Asks for add/update documentation: This type of SATD is used to request additional documentation or updates to the documentation to understand the goals and implementation of the test code. Snippet 23 provides an example, where a comment highlights the need for documentation to clarify specific behavior. [^27] https://github.com/spring-projects/spring-framework/blob/f85d5bd84a7e7bc810bb5a8179fc2fc130affc89/spring-aop/src/test/java/org/springframework/aop/aspectj/annotation/AbstractAspectJAdvisorFactoryTests.java#L435
@@
          Indicates flaky tests: This type of SATD informs developers that the test is unstable—sometimes passing and sometimes failing (i.e., flaky tests) (Eck et al., 2019). Snippet 24 provides an example of such SATD. In this instance, the comment warns that the implementation causes flaky tests. [^28] https://github.com/apache/commons-io/blob/290d72eda9152d1e11b79d48453908ff3f6b9897/src/test/java/org/apache/commons/io/FileUtilsWaitForTest.java#L42
@@
          Asks for updates of used production code in tests: This type of SATD is used to indicate outdated production code used in the test. Snippet 25 provides an example, where a comment requests updating the test to use a new conflict resolver. [^29] https://github.com/apache/maven/blob/8b094c9513efc1b9ce2d952b3b9c8eaedaf8cbf0/maven-compat/src/test/java/org/apache/maven/repository/legacy/resolver/DefaultArtifactCollectorTest.java#L155
@@
          Asks for test deletion: This type of SATD is used to report unnecessary tests. Snippet 26 provides an example, where a comment requests the deletion of a test that is no longer needed. [^30] https://github.com/apache/dubbo/blob/3609ddb2259ad223f6c0a827e36f6f8ccd38c6b2/dubbo-config/dubbo-config-api/src/test/java/org/apache/dubbo/config/MethodConfigTest.java#L111
@@
          Indicates invalid or unused tests: This type of SATD is used to inform other developers that the test is currently invalid or unused. Snippet 27 provides an example of such SATD, where the comment asks other developers to verify if the test is unused. [^31] https://github.com/argouml-tigris-org/argouml/blob/6b6db0242a40f80655cbfdddcca246afe23df20c/src/argouml-app/tests/org/argouml/model/TestAgainstUmlModel.java#L91
@@ 5.4. $RQ_{4}$: To what extent can SATD in Test Code be classified automatically?
          Next, looking into the prediction performance for subcategories, the best-performing model CodeBERT exhibits outstanding performance for predicting SATDs falling into “Test Completeness” category, (i.e., F1 score of 0.83). On the other hand, the lowest F1-score was observed in the “Failures” category with a value of 0.54. In particular, instances of the “Failures” category were frequently misclassified as “On-hold Task”. For example, we observe several cases that have a trigger condition which is often used in “On-hold Task” category such as “TODO: WriteResult isn’t returned when inserting”. [^33] https://github.com/apache/camel/blob/66736471db8ddc22e50cc31c87d34b072455b488/components/camel-mongodb/src/test/java/org/apache/camel/component/mongodb/integration/MongoDbHeaderHandlingIT.java#L71 This low performance is likely to be caused by the lack of instances (i.e., the dataset contains only 15 instances from the “Failures” category). Our future work will extend the dataset to increase the number of SATD that fall into these smaller categories.
@@ 6.1. Test Smells with Types of Test SATD
          (separated into production and test code)
@@ 6.2. Lessons Learned
          A key lesson learned is that SATD in test code is not simply a secondary concern to production code SATD; it is a distinct phenomenon with unique characteristics and developer intentions. Our identification of 20 specific types of issues, categorized into five main groups, many of which do not directly align with existing categories for production code SATD. Notably, test-specific categories like “Doubts on specific test case design” emerged from our data but were absent even from general technical debt studies (i.e., not only self-admitted technical debt). This divergence demonstrates that assuming similar characteristics for SATD across production and test code is an oversimplification.
@@
          The performance of our machine learning models varied significantly across categories, with a noticeable drop for categories with fewer instances, such as the “Failures” category, which had only 15 instances in our dataset. While the CodeBERT-based classifier achieved the highest F1-score of 0.70 overall, its performance for ”Failures” was the lowest at 0.54. This clearly demonstrates the impact of data imbalance on classification accuracy. A crucial lesson here is the need for a significantly larger and more balanced dataset to improve the robustness and accuracy of automatic SATD classification, especially for less frequent but potentially critical types of debt.

### v2

@@ 5.3.1. Production-originated Issues
Failures: This subcategory of SATD indicates the causes of test failures, which are related to either environments/settings or inputs in the production code. 15 cases fell into this category, with 11 related to environments/settings and 4 related to specific inputs. The two most frequently words are fails and enabled, which are related to test execution status. Bavota et al. (Bavota and Russo, 2016) also introduced examples of such failures, but their “Test debt” category does not contain any subcategories. Similarly, the study by Kashiwa et al. (Kashiwa et al., 2022) also presents a “Failure” category for test-related SATD, but no further categorization was given. We separate the “failure”-related SATD into two types:
  • Indicates the failure-triggering environments/settings: This type of SATD is used to notify developers that a test will fail with a specific environment or configuration. In many cases, developers comment out the assertions that fail or disable the whole test method. Snippet 7 shows an example, including an access-related failure occurring only on Linux and Windows. [^11] https://github.com/apache/commons-io/blob/290d72eda9152d1e11b79d48453908ff3f6b9897/src/test/java/org/apache/commons/io/FileUtilsTest.java#L1730 This test method is enabled to run only on Mac OS, using “@EnabledOnOs” annotation. [^12] https://junit.org/junit5/docs/5.2.0/api/org/junit/jupiter/api/condition/EnabledOnOs.html
@@
  • Indicates the failure-triggering inputs: This type of SATD notifies developers that the test will fail when the production code receives a specific input, which is illustrated in Snippet 8. In the example, the test is commented out due to a failure caused by a specific file. [^13] https://github.com/argouml-tigris-org/argouml/blob/6b6db0242a40f80655cbfdddcca246afe23df20c/src/argouml-core-model-mdr/tests/org/argouml/model/mdr/TestReadCompressedFilesAndHref.java#L73
@@
    • Indicates the existence of defects in production code: This type of SATD is used to inform developers of defects present in the production code. We identified 42 instances of this type, making it the most prevalent on-hold task. Snippet 9 shows an example of this category, where the test is currently disabled until the defect in the YARN project is fixed. [^14] https://github.com/apache/hadoop/blob/bd8b77f398f626bb7791783192ee7a5dfaeec760/hadoop-mapreduce-project/hadoop-mapreduce-client/hadoop-mapreduce-client-app/src/test/java/org/apache/hadoop/mapreduce/v2/app/TestStagingCleanup.java#L224-L227
@@
    • Indicates the unimplemented production code: This type of SATD indicates that a developer is waiting for the implementation of certain production code. An example can be seen in Snippet 10, in which a method call is commented out because a function for a property is not yet implemented. [^15] https://github.com/apache/poi/blob/ae2f0945cd2ab37260e46ab46c54b8f68a131aea/poi/src/test/java/org/apache/poi/hpsf/basic/TestWriteWellKnown.java#L230
@@
    • Specifies tasks to execute for debugging: This type of SATD informs developers of the specific task to execute for debugging. As illustrated in Snippet 11, the SATD comment requests an investigation into why the internal state is not visible during retry processing. [^16] https://github.com/apache/camel/blob/66736471db8ddc22e50cc31c87d34b072455b488/components/camel-aws/camel-aws-xray/src/test/java/org/apache/camel/component/aws/xray/ErrorTest.java#L43
@@ 5.3.2. Test-originated Issues
        Indicates workaround for impossible/difficult implementations: This type of SATD is used to inform developers about the workaround for code either impossible or difficult to implement. This is the most frequent SATD type (53 instances) regarding Test Design and Implementation. An example can be seen in Snippet 14, in which a developer attempted to retrieve the value of a property, but the current implementation of the production code prevents this due to encapsulation. Consequently, the developer resorted to using reflection to obtain the value. The comment clarifies the rationale behind this implementation choice. [^19] https://github.com/quarkusio/quarkus/blob/a91a36c533676a5d35ddedbef1717392c9191360/integration-tests/kafka-streams/src/test/java/io/quarkus/it/kafka/streams/KafkaStreamsPropertiesTest.java#L26
@@
        Indicates non-optimal way of test implementation: This type of SATD is used to highlight non-optimal implementations. Snippet 15 provides an example of such a case, where a comment points out the use of a method or approach that is suboptimal. In this instance, the SATD comment indicates that directly using DataFormatReifier is not the best practice and suggests that a better approach should be considered. [^20] https://github.com/apache/camel/blob/66736471db8ddc22e50cc31c87d34b072455b488/components/camel-fhir/camel-fhir-component/src/test/java/org/apache/camel/component/fhir/dataformat/spring/FhirDataformatConfigSpringTest.java#L90
@@
        Provides an alternative way to implement/design the test: This type of SATD is used to propose a different way to implement the test. Snippet 16 provides an example of this type of SATD, where the comment suggests using “interrupts” instead of notifyAll() to target waiting threads more effectively. [^21] https://github.com/apache/cassandra/blob/6b134265620d6b39f9771d92edd29abdfd27de6a/test/simulator/main/org/apache/cassandra/simulator/systems/InterceptingMonitors.java#L363
@@
        Doubts on specific test case design: This type of SATD raises questions about purposes, validity or effectiveness of test cases. Snippet 17 provides an example where a developer questioned the purpose of the test case and subsequently commented out the entire test method. [^22] https://github.com/apache/bookkeeper/blob/cce4b6461691466c663f2cb4d00dd4d73dd9071e/bookkeeper-server/src/test/java/org/apache/bookkeeper/bookie/datainteg/DataIntegrityCheckTest.java#L410
@@
        Doubts on test implementation: This type of SATD appears when developers are unclear about whether the implementations are optimal or not. Snippet 18 provides an example, in which the comment raises questions about the necessity of the Thread.sleep() call, suggesting that a further review is needed. [^23] https://github.com/apache/camel/blob/66736471db8ddc22e50cc31c87d34b072455b488/components/camel-github/src/test/java/org/apache/camel/component/github/consumer/PullRequestCommentConsumerTest.java#L55
@@
        Indicates the non-functional issues of tests: This type of SATD is used to inform developers about non-functional issues in the test code, such as performance, scalability, or resource usage. Snippet 19 provides an example, where the comment indicates that the test takes an excessive amount of time to execute, potentially impacting overall test efficiency. [^24] https://github.com/cloudfoundry/uaa/blob/5b74878c0297043860bb88434ee1123934acfe19/server/src/test/java/org/cloudfoundry/identity/uaa/util/CachingPasswordEncoderTest.java#L125
@@
        Test Maintenance: This subcategory of SATD is related to test code maintenance. These SATD instances indicate the need for updates in future versions, refactoring, fixing flaky tests, and adding documentation. Frequent terms such as remove and version indicate that this category captures SATD related to test code management, including cleanup and version updates. This subcategory does not exist in Bavota et al.’s work (Bavota and Russo, 2016), but “Asks for test updates in response to future/latest software versions” partially matches the “Future work” subcategory in Kashiwa et al.’s study (Kashiwa et al., 2022). We identified 48 cases in this category, with the most frequent subcategory being “Asks for test updates in response to future/latest software versions,” represented by 17 instances. The details of each subcategory are described below.
          Asks for test updates in response to future/latest software versions: This type of SATD is used to inform developers that the test needs to be updated in a future version of production code. Snippet 20 provides an example of this subcategory. The comment suggests the removal of the reference to SecurityManager in this test case when the project updates the version of the programming language Groovy, as the class is discontinued in the new version. [^25] https://github.com/apache/groovy/blob/f6221ee780bfb2f84fb197da2c13387c4e93a019/src/test/org/codehaus/groovy/reflection/SecurityTest.java#L301
          Example of SATD categorized into “Asks for test updates in response to future/latest software versions”
@@
          Asks for test code restructuring/refactoring: This type of SATD is used to notify developers that the test code needs to be restructured or refactored. Snippet 21 shows an example, where the SATD comment suggests making the contract more explicit or extracting common code to improve readability and maintainability. [^26] https://github.com/apache/flink/blob/eaffd227d853e0cdef03f1af5016e00f950930a9/flink-state-backends/flink-statebackend-changelog/src/test/java/org/apache/flink/state/changelog/ChangelogStateDiscardTest.java#L376 When looking into 9 cases, most of the SATDs request structural refactoring, including five instances of Move Method/Class, two of Extract Method, one of Rename Method, and one involving data structure changes (e.g., Introduce Parameter Object). We also found that only one case co-occurred with a test smell (specifically, “Assertion Roulette”), reinforcing our finding that Test SATD and test smells are largely independent issues.
@@
          Asks for add/update documentation: This type of SATD is used to request additional documentation or updates to the documentation to understand the goals and implementation of the test code. Snippet 22 provides an example, where a comment highlights the need for documentation to clarify specific behavior. [^27] https://github.com/spring-projects/spring-framework/blob/f85d5bd84a7e7bc810bb5a8179fc2fc130affc89/spring-aop/src/test/java/org/springframework/aop/aspectj/annotation/AbstractAspectJAdvisorFactoryTests.java#L435
@@
          Indicates flaky tests: This type of SATD informs developers that the test is unstable—sometimes passing and sometimes failing (i.e., flaky tests) (Eck et al., 2019). Snippet 23 provides an example of such SATD. In this instance, the comment warns that the implementation causes flaky tests. [^28] https://github.com/apache/commons-io/blob/290d72eda9152d1e11b79d48453908ff3f6b9897/src/test/java/org/apache/commons/io/FileUtilsWaitForTest.java#L42
@@
          Asks for updates of used production code in tests: This type of SATD is used to indicate outdated production code used in the test. Snippet 24 provides an example, where a comment requests updating the test to use a new conflict resolver. [^29] https://github.com/apache/maven/blob/8b094c9513efc1b9ce2d952b3b9c8eaedaf8cbf0/maven-compat/src/test/java/org/apache/maven/repository/legacy/resolver/DefaultArtifactCollectorTest.java#L155
@@
          Asks for test deletion: This type of SATD is used to report unnecessary tests. Snippet 25 provides an example, where a comment requests the deletion of a test that is no longer needed. [^30] https://github.com/apache/dubbo/blob/3609ddb2259ad223f6c0a827e36f6f8ccd38c6b2/dubbo-config/dubbo-config-api/src/test/java/org/apache/dubbo/config/MethodConfigTest.java#L111
@@
          Indicates invalid or unused tests: This type of SATD is used to inform other developers that the test is currently invalid or unused. Snippet 26 provides an example of such SATD, where the comment asks other developers to verify if the test is unused. [^31] https://github.com/argouml-tigris-org/argouml/blob/6b6db0242a40f80655cbfdddcca246afe23df20c/src/argouml-app/tests/org/argouml/model/TestAgainstUmlModel.java#L91
@@ 5.4. $RQ_{4}$: To what extent can SATD in Test Code be classified automatically?
          Next, looking into the prediction performance for subcategories, the best-performing model CodeBERT exhibits outstanding performance for predicting SATDs falling into “Test Completeness” category (i.e., F1 score of 0.83). On the other hand, the lowest F1-score was observed in the “Failures” category with a value of 0.54. In particular, instances of the “Failures” category were frequently misclassified as “On-hold Task”. For example, we observe several cases that have a trigger condition which is often used in “On-hold Task” category such as “TODO: WriteResult isn’t returned when inserting”. [^33] https://github.com/apache/camel/blob/66736471db8ddc22e50cc31c87d34b072455b488/components/camel-mongodb/src/test/java/org/apache/camel/component/mongodb/integration/MongoDbHeaderHandlingIT.java#L71 This low performance is likely to be caused by the lack of instances (i.e., the dataset contains only 15 instances from the “Failures” category). Our future work will extend the dataset to increase the number of SATD that fall into these smaller categories.
@@ 6.1. Test Smells with Types of Test SATD
          (separated into prod. and test code)
@@ 6.2. Lessons Learned
          A key lesson learned is that SATD in test code is not simply a secondary concern to production code SATD; it is a distinct phenomenon with unique characteristics and developer intentions. We identified 20 specific types of issues, categorized into five main groups, many of which do not directly align with existing categories for production code SATD. Notably, test-specific categories like “Doubts on specific test case design” emerged from our data but were absent even from general technical debt studies (i.e., not only self-admitted technical debt). This divergence demonstrates that assuming similar characteristics for SATD across production and test code is an oversimplification.
@@
          The performance of our machine learning models varied significantly across categories, with a noticeable drop for categories with fewer instances, such as the “Failures” category, which had only 15 instances in our dataset. While the CodeBERT-based classifier achieved the highest F1-score of 0.70 overall, its performance for “Failures” was the lowest at 0.54. This clearly demonstrates the impact of data imbalance on classification accuracy. A crucial lesson here is the need for a significantly larger and more balanced dataset to improve the robustness and accuracy of automatic SATD classification, especially for less frequent but potentially critical types of debt.